OpenStudy (muzzack):

Rewrite the rational exponent as a radical by extending the properties of integer exponents. 2 to the 7 over 8 power, all over 2 to the 1 over 4 power the eighth root of 2 to the fifth power the fifth root of 2 to the eighth power the square root of 2 to the 5 over 8 power the fourth root of 2 to the sixth power

OpenStudy (muzzack):

@musa19965

OpenStudy (muzzack):

@student_basil @Ash90 @Vonte12

OpenStudy (muzzack):

need help!!!

OpenStudy (muzzack):

\[\frac{ 2^{\frac{ 7 }{ 8 }} }{ 2 ^{\frac{ 1 }{ 4 }}}\]

OpenStudy (muzzack):

\[\sqrt[8]{2^{5}} \sqrt[5]{2^{8}} \sqrt{2^{\frac{ 5 }{ 8}}} \sqrt[4]{2^{6}}\]

OpenStudy (muzzack):

these are the choices

OpenStudy (muzzack):

@Sameerah11 @wolfe8 @wwhitlock @Warriorz13

OpenStudy (warriorz13):

The first one

OpenStudy (muzzack):

thx :) how did you know

OpenStudy (warriorz13):

openstudy.com/updates/52696bade4b040e33d2f77e9‎ here is my proof

OpenStudy (warriorz13):

Read the conversation almost at the end a person with 99 smartscore proves it

OpenStudy (muzzack):

ok thx